40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 4 12:46:27 2018, gautam, Update, ALS, Fiber ALS assay FiberALS.pdfFiberALS.svg.zip
    Reply  Thu Jan 11 14:22:40 2018, gautam, Update, ALS, Fiber ALS assay 
       Reply  Thu Jan 11 20:51:20 2018, gautam, Update, ALS, Fiber ALS assay IMG_5112.JPG
          Reply  Tue Jan 16 21:50:53 2018, gautam, Update, ALS, Fiber ALS assay 
             Reply  Thu Jan 18 00:35:00 2018, gautam, Update, ALS, Fiber ALS assay IR_ALS_20180118.pdfC2B4C1DD-6528-4067-9C13-6BD248629AD6.MOV
                Reply  Fri Jan 19 11:34:21 2018, gautam, Update, ALS, Fiber ALS assay IR_ALS_20180119.pdf
                   Reply  Fri Jan 19 23:04:11 2018, gautam, Update, ALS, Fiber ALS assay 
                      Reply  Wed Jan 24 00:33:31 2018, gautam, Update, ALS, Fiber ALS assay IMG_6866.JPG
                         Reply  Wed Jan 24 10:45:14 2018, gautam, Update, ALS, Fiber ALS assay 
                         Reply  Thu Jan 25 23:59:14 2018, gautam, Update, ALS, Fiber ALS assay 
                            Reply  Fri Jan 26 20:03:09 2018, gautam, Update, ALS, Fiber ALS assay 
                               Reply  Wed Jan 31 15:45:22 2018, gautam, Update, ALS, Fiber ALS assay IMG_6875.JPG
                                  Reply  Wed Jan 31 16:29:42 2018, gautam, Update, ALS, Modulation depths PMCmodDepth.pdfXPDH.pdf
                                  Reply  Wed Jan 31 16:33:53 2018, gautam, Update, ALS, ALS electronics at LSC rack 
                                     Reply  Wed Jan 31 22:32:11 2018, gautam, Update, ALS, ALS signal chain + power budget FiberALS_PowerBudget.pdfpreampProposed.pdf
                                        Reply  Thu Feb 1 01:24:56 2018, gautam, Update, ALS, D0902745 revamp underway PSLinterlock.png
                                           Reply  Fri Feb 2 00:26:34 2018, gautam, Update, ALS, D0902745 revamp underway 
                                              Reply  Fri Feb 2 13:16:55 2018, gautam, Update, ALS, ALS signals whitening switching ALS_whitening_switching.pdfALS_whitening_switching_works.pdf
                                                 Reply  Mon Feb 5 14:11:01 2018, gautam, Update, ALS, Huge harmonics in ALS channels ALSpowerlineHarmonics.pdf2A1B3AC4-4059-416A-AD88-8AB0431D7A55.jpeg
                                                    Reply  Mon Feb 5 22:57:28 2018, gautam, Update, ALS, Huge harmonics in ALS channels 
                                                       Reply  Tue Feb 6 11:13:26 2018, gautam, Update, ALS, Possible source of ground loop identified A68AF89C-E8A9-416D-BBD2-A1AD0A51E0B5.jpeg
                                                          Reply  Tue Feb 6 22:55:51 2018, gautam, Update, ALS, Possible source of ground loop identified 
                                                       Reply  Thu Feb 8 00:33:20 2018, gautam, Update, ALS, D990694 characterization / THD measurement plan D990694THD_trial.pdf
                                                          Reply  Thu Feb 8 01:27:16 2018, Koji, Update, ALS, D990694 characterization / THD measurement plan 
                                                          Reply  Fri Feb 9 13:37:44 2018, gautam, Update, ALS, THD measurement trial THD_trial.pdf
                                                       Reply  Thu Feb 8 12:00:09 2018, gautam, Update, ALS, D990694 is NOT differential receiving 
                                                          Reply  Thu Feb 8 13:13:14 2018, gautam, Update, ALS, D990694 pulled out IMG_5131.JPGIMG_5133.JPGD990694-B_cuttingPlan.pdf
                                                             Reply  Thu Feb 8 18:10:36 2018, gautam, Update, ALS, D990694 pulled out 
                                              Reply  Wed Feb 7 15:51:15 2018, gautam, Update, ALS, D0902745 revamp complete 6613CD37-5014-44AE-B1FE-6F6A8137DF62.jpeg
                                                 Reply  Fri Feb 16 01:34:40 2018, gautam, Update, ALS, D0902745 in-situ testing BeatMouthX_20180216.pdfIMG_5134.JPG
                                                    Reply  Tue Feb 20 23:08:27 2018, gautam, Update, ALS, D0902745 in-situ testing BeatMouthX_20180220_diffOut.pdf
                                                       Reply  Thu Feb 22 00:09:11 2018, gautam, Update, ALS, D0902745 in-situ testing daughterBoard_TF.pdfdaughterBoard_noise.pdfIMG_6906.JPG
                                                          Reply  Sun Feb 25 00:03:12 2018, gautam, Update, ALS, Daughter board prototyping schematic.pdfdaughterBoard_TF.pdfdaughterBoard_noise.pdfTEK00000.PNGdaughterBoard_noise.pdf
                                                             Reply  Mon Feb 26 20:55:56 2018, rana, Update, ALS, Daughter board prototyping 
                                                                Reply  Tue Feb 27 21:10:45 2018, gautam, Update, ALS, Daughter board testing IMG_6912.JPGIMG_6913.JPGIMG_6914.JPG
                                                                   Reply  Wed Feb 28 19:13:25 2018, gautam, Update, ALS, ADC test for differential receiving in c1lsc c1lsc_ADCtest.png
                                                                   Reply  Fri Mar 2 01:45:06 2018, gautam, Update, ALS, new look ALS electronics IMG_6916.JPG
                                                                      Reply  Mon Mar 5 17:27:34 2018, gautam, Update, ALS, new look ALS electronics - characterization BeatMouthX_20180305_diffOut.pdf
                                                                         Reply  Thu Mar 8 00:40:25 2018, gautam, Update, ALS, new look ALS electronics - characterization BeatMouthX_20180305_diffOut.pdf
                                                                            Reply  Thu Mar 8 19:38:37 2018, gautam, Update, ALS, digital unwhitening of daughter board TFfit.pdf
                                                                               Reply  Thu Mar 8 23:50:27 2018, gautam, Update, ALS, First look at new ALS electronics BeatMouth_OOL.pdf
                                                                                  Reply  Fri Mar 9 01:07:01 2018, gautam, Update, ALS, First look at new ALS electronics ALS_vs_POXnoise.pdf
                                                                                     Reply  Mon Mar 12 22:08:31 2018, gautam, Update, ALS, Noisy POX NoisyPOX_20180312.pdf
                                                                                        Reply  Mon Mar 12 23:57:31 2018, gautam, Update, ALS, Noisy POX NoisyPOXandPOY_20180312.pdfnoisyPOX_Jan2018.pdf
                                                                                           Reply  Mon Mar 19 15:02:29 2018, gautam, Update, ALS, Noisy MC sensing 
                                        Reply  Thu Feb 1 15:31:12 2018, gautam, Update, ALS, ALS signal chain + power budget 
                                        Reply  Wed Feb 7 10:16:26 2018, gautam, Update, ALS, ALS signal chain + power budget 
                      Reply  Thu Jan 25 13:18:41 2018, gautam, Update, ALS, Fiber ALS assay preampProposed.pdf
Message ID: 13680     Entry time: Mon Mar 12 23:57:31 2018     In reply to: 13679     Reply to this: 13688
Author: gautam 
Type: Update 
Category: ALS 
Subject: Noisy POX 

[kevin, gautam]

Kevin suggested I shouldn't be so lazy and test the POY spectrum as well. So we moved the timing card back to c1iscey, went through the usual dance of vertex machine reboots, and then got both single arm locks going. Attached spectrum shows that both POX and POY are noisy. I'm not sure what has changed that could cause this effect. The fact that both POX and POY appear uniformly bad, but that there is no coherence with MC_F, suggests to me that perhaps this has something to do with the work I did with Koji w.r.t. the power situation at the LSC rack. But we just checked that

  1. All the demod board front panel LED indicators are green.
  2. Marconi and all RF amplifier boxes are on (but we didn't actually measure any RF power levels yet tonight).
  3. We checked the KEPCO power supplies in the little cabinet along the Yarm, and all of them are reporting the correct voltages/currents as per Steve's (recently updated) labels.
  4. Checked the expansion chassis at the LSC rack for any red lights, there were none.

Another observation we made: note the huge bump around 70Hz in both arm control signals. We don't know what the cause of this is. But we occassionally noticed harmonics of this (i.e. 140, 210 Hz etc) appear in the control signal spectra, and they would grow with time - eventually, the X arm would lose lock (though the Y arm stayed locked).

I'm short on ideas for now so we will continue debugging tomorrow.


Unrelated to this work: Kevin reminded me that the high-pitched whine from the CRT TVs in the control room (which is apparently due to the flyback transformer) is DEAFENING. It's curious that the "chirp" to the eventual 15kHz whine is in opposite directions for the QUAD CRTs and the single display ones. Should be a Ph6 experiment maybe.


Update 2:30pm Mar 13: The furthest back I seem to be able to go in time with Frames is ~Jan 20 2018. Looking for a time when the arms were locked from back then, it seems like whatever is responsible for a noisy POX and POY was already a problem back in January. See Attachment #2. So it appears that the recent work at 1Y2 is not to blame...

Attachment 1: NoisyPOXandPOY_20180312.pdf  74 kB  Uploaded Tue Mar 13 01:04:45 2018  | Hide | Hide all
NoisyPOXandPOY_20180312.pdf
Attachment 2: noisyPOX_Jan2018.pdf  9 kB  Uploaded Tue Mar 13 15:21:56 2018  | Hide | Hide all
noisyPOX_Jan2018.pdf
ELOG V3.1.3-